# Enhancing Physics-Informed Neural Networks Through Feature Engineering

### v2

@@ B.4 Heat
$(\mathbf{x},t)\in\Omega\times T=[0,1]^{2}\times[0,5],$
@@
The analytical solution is not readily available for every condition, but for this specific study, we use the dataset for the 2D heat task from PINNacle (see (Hao et al., 2023)).
@@ B.6 Diffusion
This equation is usually available in PDE benchmarks as ”reaction-diffusion”, which is a different equation. In our study, we wanted to include this PDE to compare results with to the previous similar studies such as (Zeng et al., 2024b), so since this PDE as ”diffusion” only was not available in online benchmarks like PDEBench, and the analytical solution is available, it is simply implemented in our study the way it is implemented in the previous similar studies.
@@ B.7 Allen-Cahn
The analytical solution to this PDE can be derived under certain conditions. In this study, we used Allen-Cahn’s implementation from PredictiveIntelligenceLab/jaxpi; see (Wang et al., 2024).
@@ C.1.1 Ablation Study on Activation Functions
To validate our selection of the tanh activation function for SAFE-NET’s hidden layer, we conduct an ablation study comparing five commonly used activation functions in PINN architectures: tanh, sine (as used in SIREN (Sitzmann et al., 2020)), ReLU, GELU, and Swish. Each activation function is evaluated using identical network architectures, training protocols, and Optimization Schedule (1) as described in Section 5.2.
@@ C.1.2 Ablation Study on the Number of Features
As evident from Figure 9, the number of features in SAFE-NET directly impacts both the representational capacity and computational complexity of the network. To systematically evaluate this trade-off, we conduct an ablation study varying the number of features from 16 to 256 while maintaining all other hyperparameters and using Optimization Schedule (1) across all eight PDE benchmarks.
@@ C.1.3 Fourier Coefficient and Frequencies Initializations
Frequency Initialization Strategies: We compare three different approaches for initializing the trainable frequencies $\omega_{x}^{(i)}$ and $\omega_{t}^{(i)}$:
@@
    Harmonic Frequencies (Ours): $\omega_{x}^{(i)}=\omega_{t}^{(i)}=i\pi$ for $i=1,2,\ldots,N$
@@
    Random Gaussian: $\omega_{x}^{(i)},\omega_{t}^{(i)}\sim\mathcal{N}(0,\pi^{2})$
@@
    Uniform Random: $\omega_{x}^{(i)},\omega_{t}^{(i)}\sim\mathcal{U}(0,2\pi)$
Coefficient Initialization Strategies: We evaluate four initialization schemes for the Fourier coefficients $\text{coeff}_{j}^{(i)}$:
@@
    Random Gaussian: $\text{coeff}_{j}^{(i)}\sim\mathcal{N}(0,1)$
@@
    Random Uniform: $\text{coeff}_{j}^{(i)}\sim\mathcal{U}(0,1)$
@@
    Xavier Initialization: $\text{coeff}_{j}^{(i)}\sim\mathcal{U}(-\sqrt{6/n},\sqrt{6/n})$ where $n$ is the input dimension
@@ C.1.4 Architecture Depth Ablation
Figure 19 presents a pattern: while increasing from 1 to 4 hidden layers provides modest improvements (average $\times1$ improvement from $1\rightarrow 2$ layers and additional $\times1$ from $2\rightarrow 4$ layers), further depth increases show minimal performance changes compared to 4-layer networks, indicating a performance plateau.
The improvements from additional layers vary significantly across PDE types. Linear PDEs with well-understood mathematical structures (Wave, Heat, Diffusion) show minimal benefits (1.031.05×1.03-1.05$\times$1.03 - 1.05 × improvement), while nonlinear PDEs with complex dynamics demonstrate more substantial gains: Burgers equation achieves 1.5×1.5$\times$1.5 × improvement due to its sharp discontinuities, Allen-Cahn shows 1.2×1.2$\times$1.2 × improvement, and Navier-Stokes exhibits 1.2×1.2$\times$1.2 × improvement for complex flow physics. This differential behavior supports our hypothesis that feature engineering is most effective for PDEs with known mathematical structures.
@@ C.1.5 Feature Normalization Ablation
Figure 20shows that linear PDEs with smooth solutions (Wave, Heat, Diffusion, Reaction) show stable but significantly degraded performance (410×4-10$\times$4 - 10 × worse errors) without normalization, while nonlinear PDEs with complex dynamics (Convection, Allen-Cahn, Burgers, Navier-Stokes) exhibit complete L-BFGS optimization failure.
@@ C.2 Complete Results for Figure 1
C.3 Spectral Density Computation
@@
C.3.1 Spectral Density Computation Methodology
@@
Computing the full eigendecomposition of the Hessian is computationally prohibitive for neural networks with thousands of parameters. Following the methodology established in (Yao et al., 2020) and (Golub & Meurant, 2009), we employ Stochastic Lanczos Quadrature (SLQ) to efficiently approximate the spectral density. SLQ requires only Hessian-vector products, which can be computed efficiently using automatic differentiation.
@@
For our implementation, we utilize the PyHessian library (Yao et al., 2020), which provides efficient GPU-accelerated computation of spectral densities for neural networks. We use $n_{v}=100$ stochastic samples and a Lanczos iteration count of 200 to ensure accurate spectral density estimation across all experiments.
C.3.2 Experimental Setup for Spectral Analysis
@@
To ensure fair comparison, SAFE-NET and all baseline methods maintain identical parameter counts where possible. SAFE-NET uses 128 features as in the main experiments, while baseline methods use their standard configurations. Network weights are initialized using Xavier initialization (Glorot & Bengio, 2010) across all methods.
@@
C.3.3 Results and Analysis
Figure 10 demonstrates the spectral density comparison between SAFE-NET and standard PINN for the wave equation, showing dramatic improvements in conditioning. The comprehensive results across all PDEs are presented in Tables 20 and 21. Also, Figures 2125 are provided as visualization examples.
@@ C.3.3 Results and Analysis
Reduction Factor (×$\times 10^{3}$) (×$\times 10^{1}$) (×$\times 10^{2}$) (×$\times 10^{1}$) (×$\times 10^{2}$) (×$\times 10^{2}$) (×$\times 10^{3}$)
PINN 1.2×$\times 10^{4}$ 2.0×$\times 10^{1}$ 6.6×$\times 10^{2}$ 1.1×$\times 10^{2}$ $\times 10^{2}$ 1.6×$\times 10^{2}$ 1.5×$\times 10^{4}$
FLS-PINN 8.8×$\times 10^{3}$ 1.9×$\times 10^{1}$ 4.9×$\times 10^{2}$ 9.5×$\times 10^{1}$ 1.9×$\times 10^{2}$ 1.5×$\times 10^{2}$ 1.3×$\times 10^{4}$
W-PINN 7.9×$\times 10^{3}$ 1.8×$\times 10^{1}$ 4.4×$\times 10^{2}$ 8.8×$\times 10^{1}$ 1.7×$\times 10^{2}$ 1.3×$\times 10^{2}$ 1.1×$\times 10^{4}$
RBA-PINN 3.8×$\times 10^{2}$ 8.3×$\times 10^{0}$ 1.9×$\times 10^{2}$ 3.9×$\times 10^{1}$ 1.1×$\times 10^{2}$ 7.8×$\times 10^{1}$ 1.4×$\times 10^{3}$
RFF 1.6×$\times 10^{3}$ 1.2×$\times 10^{1}$ 3.1×$\times 10^{2}$ 5.9×$\times 10^{1}$ 1.4×$\times 10^{2}$ 1.1×$\times 10^{2}$ 3.5×$\times 10^{3}$
RBF 6.0×$\times 10^{2}$ 1.1×$\times 10^{1}$ 2.5×$\times 10^{2}$ 5.1×$\times 10^{1}$ 1.2×$\times 10^{2}$ 1.0×$\times 10^{2}$ 2.7×$\times 10^{3}$
RBF-P 3.1×$\times 10^{1}$ 8.0×$\times 10^{0}$ 1.7×$\times 10^{2}$ 3.8×$\times 10^{1}$ 1.1×$\times 10^{2}$ 8.8×$\times 10^{1}$ 1.9×$\times 10^{3}$
@@
Figures, 10(a), 21(a), 22(a), 23(a), 24(a), and 25(a) display spectral density plots at early training stages (3,000 iterations). Even at this early phase, SAFE-NET demonstrates improved conditioning compared to standard PINNs. The eigenvalue distribution is more concentrated around moderate values, with fewer extremely large eigenvalues that typically cause ill-conditioning. This suggests that SAFE-NET’s feature engineering provides inherently better conditioning from initialization, facilitating more stable gradient-based optimization from the onset of training.
@@
The spectral density plots at the end of training (Figures 10(b), 21(b), 22(b), 23(b), 24(b), and 25(b) ) reveal even more dramatic conditioning improvements. Specifically for SAFE-NET and PINN:
@@
Figure 21: Spectral density plots at the beginning and end of training for the wave PDE with RFF and RBF. Comapre with Figure 10 demonstrating spectral density for PINN and SAFE-NET

### v3

@@ B.4 Heat
$(x,y,t)\in\Omega\times T=[0,1]^{2}\times[0,5],$
@@
The analytical solution is not readily available for every condition, but for this specific study, we use the dataset for the 2D heat task from PINNacle (see hao2023pinnacle).
@@ B.6 Diffusion
This equation is usually available in PDE benchmarks as "reaction-diffusion", which is a different equation. In our study, we wanted to include this PDE to compare results with to the previous similar studies such as zeng2024rbfpinnnonfourierpositionalembedding, so since this PDE as "diffusion" only was not available in online benchmarks like PDEBench, and the analytical solution is available, it is simply implemented in our study the way it is implemented in the previous similar studies.
@@ B.7 Allen-Cahn
The analytical solution to this PDE can be derived under certain conditions. In this study, we used Allen-Cahn’s implementation from PredictiveIntelligenceLab/jaxpi; see wang2024piratenets.
@@ C.1.1 Ablation Study on Activation Functions
To validate our selection of the tanh activation function for SAFE-NET’s hidden layer, we conduct an ablation study comparing five commonly used activation functions in PINN architectures: tanh, sine (as used in SIREN sitzmann2020implicit), ReLU, GELU, and Swish. Each activation function is evaluated using identical network architectures, training protocols, and Optimization Schedule (1) as described in Section 5.2.
@@ C.1.2 Ablation Study on the Number of Features
As evident from Figure˜9, the number of features in SAFE-NET directly impacts both the representational capacity and computational complexity of the network. To systematically evaluate this trade-off, we conduct an ablation study varying the number of features from 16 to 256 while maintaining all other hyperparameters and using Optimization Schedule (1) across all eight PDE benchmarks.
@@ C.1.3 Fourier Coefficient and Frequencies Initializations
Frequency Initialization Strategies: We compare three different approaches for initializing the trainable frequencies $\omega_{x}^{(\ell)}$ and $\lambda_{t}^{(\ell)}$:
@@
    Harmonic Frequencies (Ours): $\omega_{x}^{(\ell)}=\lambda_{t}^{(\ell)}=i\pi$ for $\ell=1,2,\ldots,N$
@@
    Random Gaussian: $\omega_{x}^{(\ell)},\lambda_{t}^{(\ell)}\sim\mathcal{N}(0,\pi^{2})$
@@
    Uniform Random: $\omega_{x}^{(\ell)},\lambda_{t}^{(\ell)}\sim\mathcal{U}(0,2\pi)$
Coefficient Initialization Strategies: We evaluate four initialization schemes for the Fourier coefficients $\text{coeff}_{j}^{(\ell)}$:
@@
    Random Gaussian: $\text{coeff}_{j}^{(\ell)}\sim\mathcal{N}(0,1)$
@@
    Random Uniform: $\text{coeff}_{j}^{(\ell)}\sim\mathcal{U}(0,1)$
@@
    Xavier Initialization: $\text{coeff}_{j}^{(\ell)}\sim\mathcal{U}(-\sqrt{6/n},\sqrt{6/n})$ where $n$ is the input dimension
@@ C.1.4 Architecture Depth Ablation
Figure 19 presents a pattern: while increasing from 1 to 4 hidden layers provides modest improvements (average $1.1\times$ improvement from $1\rightarrow 2$ layers and additional $1.1\times$ from $2\rightarrow 4$ layers), further depth increases show minimal performance changes compared to 4-layer networks, indicating a performance plateau.
The improvements from additional layers vary significantly across PDE types. Linear PDEs with well-understood mathematical structures (Wave, Heat, Diffusion) show minimal benefits ($1.03-1.05\times$ improvement), while nonlinear PDEs with complex dynamics demonstrate more substantial gains: Burgers equation achieves $1.5\times$ improvement due to its sharp discontinuities, Allen-Cahn shows $1.2\times$ improvement, and Navier-Stokes exhibits $1.2\times$ improvement for complex flow physics. This differential behavior supports our hypothesis that feature engineering is most effective for PDEs with known mathematical structures.
@@ C.1.5 Feature Normalization Ablation
Figure 20 shows that linear PDEs with smooth solutions (Wave, Heat, Diffusion, Reaction) show stable but significantly degraded performance ($4-10\times$ worse errors) without normalization, while nonlinear PDEs with complex dynamics (Convection, Allen-Cahn, Burgers, Navier-Stokes) exhibit complete L-BFGS optimization failure.
C.1.6 The role of Domain Knowledge Features (DKF)
We compare the following four variants;
  • PINN (standard baseline PINN)
  • SAFE-NET (default SAFE-NET with DKF)
  • PINN ++ DKF (baseline PINN with DKF added)
  • SAFE-NET - DKF (SAFE-NET without the DKF component)
on the same eight PDE benchmarks used in our experiments. Figure 21 demonstrates the results.
Figure 21: Relative $\ell_{2}$ error across all eight PDE benchmarks for PINN and SAFE-NET with and without DKF
Training protocol.
The loss weights, number of features, network widths, sampling strategy, and all other hyperparameters follow the default experimental setup used throughout the paper (See Appendix A). Each result is averaged over five initializations with Optimization Schedule (2).
Experimental results suggest that DKF provides a sizable boost when the initial–boundary conditions might strongly influence the solution (such as Wave), and only marginal boosts when dynamics dominate (such as Burgers). Adding DKF to the baseline PINN improves the final L2RE by about 8% on average. Across all eight PDEs, SAFE-NET’s adaptive Fourier features alone reduce error by an average of 87.2% relative to the baseline PINN. Incorporating DKF into SAFE-NET yields an additional average reduction of 15.6%, with the largest gain on the Wave equation (49.8%) and smaller gains on Burgers, Navier–Stokes, Convection, and Diffusion as seen in Figure 21. These observations indicate that SAFE-NET’s adaptive Fourier features already capture the dominant dynamics of the latter equations.
@@ C.2 Complete Results for Figure 1
C.3 Additional Experiments with Non-homogeneous Boundary Conditions
In this section, we add a test case for non-homogeneous boundary conditions using a non-homogeneous boundary variant of the heat equation to test our proposed framework. Table 20 reports the final relative $\ell_{2}$ errors (L2RE).
PDE Description.
We use the following non-homogeneous heat task with mixed Dirichlet/Robin-style boundary conditions:
$u_{t}-u_{xx}-f(x,t)=0\qquad(x,t)\in[0,1]\times[0,1].$
Initial condition
$u(x,0)=\sin(2\pi x)+0.5\,x+1.$
Boundary conditions
$u(0,t)=1,\qquad u_{x}(1,t)+h\,u(1,t)=g(t),\qquad h=4.$
Forcing term
$f(x,t)=\;$ $4\pi^{2}\sin(2\pi x)\cos(\pi t)-\pi\sin(2\pi x)\sin(\pi t)$
$+\frac{3}{5}(6.3\pi)^{2}\sin(6.3\pi x)\sin(3\pi t)+\frac{3}{5}\,3\pi\,\sin(6.3\pi x)\cos(3\pi t).$
Robin data
$g(t)=\;$ $2\pi\cos(\pi t)+\frac{189\pi}{50}\cos(6.3\pi)\sin(3\pi t)+4\left[\tfrac{3}{5}\sin(6.3\pi)\sin(3\pi t)+1.5\right]+0.5$
$=2\pi\cos(\pi t)+3.78\pi\cos(6.3\pi)\sin(3\pi t)+2.4\sin(6.3\pi)\sin(3\pi t)+6.5$
Analytical solution
$u(x,t)=\sin(2\pi x)\cos(\pi t)+\frac{3}{5}\sin(6.3\pi x)\sin(3\pi t)+0.5\,x+1.$
Training protocol.
The loss weights, number of features, network widths, sampling strategy, and all other hyperparameters follow the default experimental setup used throughout the paper (See Appendix A). Each result is averaged over five initializations with Optimization Schedule (2).
Table 20: Non-homogeneous heat. best L2RE results in bold.
Method L2RE
PINN $2.92\mathrm{e}{-3}$
FLS-PINN $3.08\mathrm{e}{-3}$
W-PINN $7.01\mathrm{e}{-3}$
RFF $5.12\mathrm{e}{-3}$
RBF $9.28\mathrm{e}{-4}$
RBF-P $3.06\mathrm{e}{-4}$
RBA-PINN $3.22\mathrm{e}{-4}$
SAFE-NET $\mathbf{2.11\mathrm{e}{-4}}$
C.4 Spectral Density Computation
@@
C.4.1 Spectral Density Computation Methodology
@@
Computing the full eigendecomposition of the Hessian is computationally prohibitive for neural networks with thousands of parameters. Following the methodology established in yao2020pyhessian and golub2009matrices, we employ Stochastic Lanczos Quadrature (SLQ) to efficiently approximate the spectral density. SLQ requires only Hessian-vector products, which can be computed efficiently using automatic differentiation.
@@
For our implementation, we utilize the PyHessian library yao2020pyhessian, which provides efficient GPU-accelerated computation of spectral densities for neural networks. We use $n_{v}=100$ stochastic samples and a Lanczos iteration count of 200 to ensure accurate spectral density estimation across all experiments.
C.4.2 Experimental Setup for Spectral Analysis
@@
To ensure fair comparison, SAFE-NET and all baseline methods maintain identical parameter counts where possible. SAFE-NET uses 128 features as in the main experiments, while baseline methods use their standard configurations. Network weights are initialized using Xavier initialization glorot2010understanding across all methods.
@@
C.4.3 Results and Analysis
Figure 10 demonstrates the spectral density comparison between SAFE-NET and standard PINN for the wave equation, showing dramatic improvements in conditioning. The comprehensive results across all PDEs are presented in Tables 21 and 22. Also, Figures 2226 are provided as visualization examples.
@@ C.4.3 Results and Analysis
Reduction Factor ($\times 10^{3}$) ($\times 10^{1}$) ($\times 10^{2}$) ($\times 10^{1}$) ($\times 10^{2}$) ($\times 10^{2}$) ($\times 10^{3}$)
[…]
@@
Figures, 10(a), 22(a), 23(a), 24(a), 25(a), and 26(a) display spectral density plots at early training stages (3,000 iterations). Even at this early phase, SAFE-NET demonstrates improved conditioning compared to standard PINNs. The eigenvalue distribution is more concentrated around moderate values, with fewer extremely large eigenvalues that typically cause ill-conditioning. This suggests that SAFE-NET’s feature engineering provides inherently better conditioning from initialization, facilitating more stable gradient-based optimization from the onset of training.
@@
The spectral density plots at the end of training (Figures 10(b), 22(b), 23(b), 24(b), 25(b), and 26(b) ) reveal even more dramatic conditioning improvements. Specifically for SAFE-NET and PINN:
@@
Figure 22: Spectral density plots at the beginning and end of training for the wave PDE with RFF and RBF. Compare with Figure 10 demonstrating spectral density for PINN and SAFE-NET
@@
Appendix D A Comparison with Large-scale Architectures on Shared Benchmarks
We include the following reported numbers verbatim from the original papers of FNO li2020fno, PirateNets wang2024piratenets, PINNsFormer zhao2023pinnsformer as examples of large complex models only for the sake of completeness and to give readers an at-a-glance sense of scale (parameters, memory, time/epoch) and accuracy on overlapping PDE families, however, as each method utilizes distinct settings, we do not provide a direct ranking between them. Note that FNO is an operator-learning method that learns mappings between function spaces (e.g., initial condition $\mapsto$ solution), whereas SAFE-NET (and the PINN-style baselines considered in our main text) solve a single PDE instance with specified IC/BC. Formulations, metrics, and datasets therefore differ. The following is intended only to document the resource scale and the published accuracy on broadly overlapping PDE families.
FNO Reports results for 1D Burgers and 2D Navier–Stokes (space–time operator learning). Hardware noted by the authors: single NVIDIA V100 16 GB.
Table 23: FNO on 1D Burgers (relative $\ell_{2}$ error at different spatial resolutions $s$).
Method s=256s{=}256 $512$ $1024$ $2048$ $4096$ $8192$
FNO 0.0149 0.0158 0.0160 0.0146 0.0142 0.0139
Notes. Table reproduced from the paper; parameters, GPU memory, and time/epoch were not reported for the Burgers experiment. See Table 24 for Navier–Stokes resource numbers as reported by the authors.
Table 24: FNO on 2D Navier–Stokes (relative $\ell_{2}$ error over different viscosities $\nu$ and dataset sizes NN; per‑epoch time reported by the authors).
Method Params Time/epoch $\nu{=}10^{-3}$, T=50T{=}50, N=1000N{=}1000 $\nu{=}10^{-4}$, T=30T{=}30, N=1000N{=}1000 $\nu{=}10^{-4}$, T=30T{=}30, N=10000N{=}10000 $\nu{=}10^{-5}$, T=20T{=}20, N=1000N{=}1000
FNO-3D 6,558,537 38.99 s 0.0086 0.1918 0.0820 0.1893
FNO-2D 414,517 127.80 s 0.0128 0.1559 0.0973 0.1556
U -Net 24,950,491 48.67 s 0.0245 0.2051 0.1190 0.1982
TF -Net 7,451,724 47.21 s 0.0225 0.2253 0.1168 0.2268
ResNet 266,641 78.47 s 0.0701 0.2871 0.2311 0.2753
Notes. Reported at $64{\times}64$ spatial resolution; FNO-3D convolves in space–time while FNO-2D uses 2D convolutions with an RNN in time.
PirateNets has PINN backbone with physics-informed residual adaptive blocks. The paper emphasizes accuracy comparisons and ablations; it does not tabulate parameter counts, GPU memory, or wall-clock per epoch. Below we list the state-of-the-art test errors the authors report.
Table 25: PirateNets: reported relative $\ell_{2}$ test errors across PDEs (paper’s Table 1).
Benchmark Error (PirateNet) Params GPU Mem Time/epoch
Allen–Cahn (1D) $2.24{\times}10^{-5}$
Korteweg–De Vries (1D) $4.27{\times}10^{-4}$
Grey–Scott (2D) $3.61{\times}10^{-3}$
Ginzburg–Landau (2D) $1.49{\times}10^{-2}$
Lid-driven cavity (2D) $4.21{\times}10^{-2}$
Notes. Architecture details (e.g., depth/width) and training pipelines are provided, but resource metrics are not tabulated.
PINNsFormer is a transformer-style PINN variant. The authors report parameter counts and training overhead (V100), and test errors on overlapping 1D PDEs.
Table 26: PINNsFormer: model size and training overhead (Appendix Table 4–5 in the paper).
Model Params GPU Mem (MiB) Time/epoch (s)
PINNsFormer (pseudo-seq. length k=5k{=}5) 454,000 2,827 2.34
Notes. Reported on a single NVIDIA Tesla V100; overheads shown for k=5k{=}5.
Table 27: PINNsFormer: reported test errors on 1D PDEs used widely in PINN literature.
PDE (dimension) Metric (paper) Error Params Time/epoch / GPU Mem
Convection (1D) rRMSE ($\approx$ rel. $\ell_{2}$) 0.027 454k 2.34 s / 2,827 MiB
Reaction (1D) rRMSE ($\approx$ rel. $\ell_{2}$) 0.030 454k 2.34 s / 2,827 MiB
Notes. Errors are taken directly from the paper’s main results tables; rRMSE is the paper’s standard relative $\ell_{2}$ metric. The reaction/convection formulations and sampling follow the setups specified in zhao2023pinnsformer.
Appendix E A Discussion on Relevant Ideas in Different Structures
In this section we discuss three different frameworks that leverage Fourier structure for PDEs in different ways; SAFE-NET, FNO li2020fno, and EPGP härkönen2023gaussianprocesspriorssystems. Although all three exploit Fourier structure, they do so for different learning objectives and with complementary inductive biases.
Problem formulation and learning objective.
SAFE-NET is a physics-informed solver for a single PDE instance: we fit $u_{\theta}$ by minimizing a residual-based PINN loss with collocation points for the PDE, BCs and (if present) ICs; no paired input-output simulation dataset is required. FNO li2020fno, by contrast, learns a resolution-invariant solution operator $G^{\dagger}\!:a\mapsto u$ from many paired fields $\{(a^{(i)},u^{(i)})\}_{i=1}^{N}$; once trained it solves new instances in one forward pass and even on unseen grids (zero-shot super-resolution). EPGP from härkönen2023gaussianprocesspriorssystems is probabilistic: it places a GP prior whose sample paths satisfy a linear, constant-coefficient PDE exactly; conditioning on scattered observations (values and/or derivatives) yields a posterior that remains within the PDE solution set.
Spectral parameterization.
All three approaches leverage Fourier structure, but in different ways.
  • SAFE-NET uses a finite, trainable tensor-product Fourier feature map in space (and time), with sine and cosine cross terms and learned amplitudes. Both the harmonically-initialized frequencies and amplitudes are optimized jointly with the network via the physics loss. This couples spectral adaptivity to the governing equation and initial-boundary condition information embedded in the loss.
  • FNO parameterizes convolutions in Fourier space: the FFT of the feature field is multiplied by learnable complex matrices on a fixed set of retained modes and transformed back by an inverse FFT. In other words, each Fourier layer applies $v_{t+1}(x)=\mathcal{F}^{-1}\!\bigl(R\cdot\mathcal{F}v_{t}\bigr)(x)+Wv_{t}(x),$ where $R$ is a learnable complex tensor restricted to low modes and $W$ a local linear map; frequencies themselves remain fixed and are not learned; instead, a (problem-dependent) truncation $k_{\max}$ selects the modes that are updated, while higher modes are filtered out li2020fno. This yields global, resolution-invariant layers well suited to operator learning.
  • EPGP in härkönen2023gaussianprocesspriorssystems expresses solutions through the Ehrenpreis–Palamodov representation (a spectral integral over the PDE’s characteristic variety) as the following kernel
    $k_{\text{EPGP}}(x,x^{\prime})=\!\!\int_{\sqrt{-1}\mathbb{R}^{d}}\!\!\Psi(x,z)\,\Psi^{\!*}(x^{\prime},z)\,e^{-\|z\|^{2}/2}\,{\rm d}z,$ (46)
    with $\Psi(x,z)=\sum_{j,z\in S_{z^{\prime}}}D_{j}(x,z)\,\exp\langle x,z\rangle$. A Monte-Carlo variant (S-EPGP) replaces the integral by a finite sum of trainable plane waves, so the relevant spectral locations are learned within the GP kernel.
Data/physics requirements and scope.
SAFE-NET relies on the PDE residual (and BC/IC) and targets low-to-moderate‑dimensional PDEs (including smooth and mildly non-smooth regimes), where Fourier bases provide a good inductive bias. FNO excels when sufficiently many input–output pairs of fields are available; its strength is generalization across instances and grids rather than solving a single instance from physics alone. EPGP is model-based and probabilistic: it assumes linear PDEs with constant coefficients and incorporates data at arbitrary locations (values and/or derivatives).
Boundary/initial conditions and domain assumptions.
SAFE-NET feeds BC/IC conditions, if available, directly as domain features (e.g., sine series that satisfy homogeneous Dirichlet BCs) and trains end-to-end with the PINN loss. FNO typically assumes periodic structure within each Fourier layer (the convolutional kernel is defined for periodic functions), although practical variants can handle nonperiodic domains with padding or localized corrections. EPGP does not require prescribing BCs/ICs to define the prior; when BC/IC observations are available, they are simply assimilated as GP conditioning data, and the posterior remains in the PDE solution set.
Computational aspects.
Each of these architectures solve different problems so they should not be compared directly based on computational cost. However, for the sake of completion, we provide computational footprint details for each method. SAFE-NET uses a single hidden layer and a small trainable feature module, giving low parameter counts and fast epochs suitable for residual-based training. FNO stacks multiple FFT-based layers; training scales as $\mathcal{O}(Nn\log n)$ but inference on new grids is extremely fast. EPGP shifts cost to kernel construction $\mathcal{O}(n^{3})$ (or $\mathcal{O}(nm^{2})$ for its sparse variant) while providing calibrated uncertainty. With sparse/Monte–Carlo kernels, the dominant cost is one-time covariance construction.
Complementarity.
FNO works well when abundant multi-instance data are available, offering orders-of-magnitude speed-ups at deployment and results generalizable within the parametrized family of PDEs. SAFE-NET is best when the PDE (best for PDEs with moderately smooth behavior) is known but training data are scarce, computational budget is limited, or when we seek both accuracy and efficiency together. EPGP delivers principled uncertainty for linear PDEs and fuses scattered measurements seamlessly.